You buy a computer directly from the manufacturer for R7332, 00 and agree to repay it in equal installments over three years at the end of each month starting one month from now. The interest rate is 10,7% per year, compounded monthly. How much interest will you pay in total?


1
Expert's answer
2021-05-07T09:01:01-0400

Given the following:

Future value (FV)= R7332

Rate (r)= 10.7% or 0.107

Time (n)= 3 years

The installmental parent is given by

"PV=\frac{FV}{[\frac{(1+r)^n-1}{r}]}"


"PV=\frac{7332}{[\frac{(1+0.107)^3-1}{0.107}]}"


"PV=R2200.18"


The picture above shows the payment schedule.


The total interest due in the payment is R731.44 (235.42+496.02)